OpenStudy (anonymous):

Find the general solution to y*dy/dx=4^x

OpenStudy (anonymous):

Ok, so we have $$ y \cdot \frac{dy}{dx} = 4^x \\ y \cdot dy = 4^x \cdot dx \\ \int y \cdot dy = \int 4^x \cdot dx\\ \frac{y^2}{2} = \frac{4^x}{\ln(4)} + C \\ y^2 = \frac{2 \cdot 4^x}{\ln(4)} + C \\ y = \pm\sqrt{\frac{2 \cdot 4^x}{\ln(4)} + C} $$

OpenStudy (anonymous):

Thanks, I'm embarrassed I couldn't get this. I was just confused as to why Microsoft Math wasn't giving me that answer, but I think it's an issue with the program.

OpenStudy (anonymous):

Lol don't worry that's ok =) You can use wolfram alpha for this http://www.wolframalpha.com/input/?i=y*dy%2Fdx%3D4%5Ex